OpenStudy (anonymous):

While simplifying some math work, Peter wrote on his paper that x3 • x3 • x3 • x3 equaled x3+ 3 + 3 +3 . Did Peter simplify his work correctly and completely to a final answer? Would Peter’s work be the same if he were to simplify x3 + x3 + x3 + x3? Simplify the given expression to rational exponent form, justify each step by identifying the properties of rational exponents used. All work must be shown. 1 over the cube root of the quantity of x to the negative sixth power Simplify the given expression to radical form, justify each step by identifying the properties of rat

OpenStudy (anonymous):

yes peter's work is right in the first case

OpenStudy (atlas):

Hey your question would be much clearer if you use 'equation' tool. Right now your question can have multiple meanings

zepdrix (zepdrix):

\[\Large x^3\cdot x^3\cdot x^3\cdot x^3\quad=\quad x^{3+3+3+3}\] `Did Peter simplify his work correctly`, yes. `and completely to a final answer?` No, it can be simplified further.

OpenStudy (anonymous):

since 3+3+3+3=12 lol